OpenStudy (anonymous):

Which equation can be used to solve the following word problem? Emilio has 6 more nickels than dimes, and the total value of his coins is $2.10. How many nickels does he have? Let n represent the number of nickels.

OpenStudy (anonymous):

I'm dying.

OpenStudy (anonymous):

@beccaboo333 @charlotte123 @jim_thompson5910 @BlueEyedCountryGirl

OpenStudy (anonymous):

A. 0.10(n + 6) + 0.05n = 2.10 B. 0.10n + 0.05n + 6 = 2.10 C. 0.10n + 0.05(n + 6) = 2.10 D. 0.10(n – 6) + 0.05n = 2.10

OpenStudy (shamil98):

dimes = n – 6

OpenStudy (anonymous):

@beccaboo333 Thanks anyways :)

OpenStudy (shamil98):

0.05n + 0.10(n-6) = 2.10

OpenStudy (anonymous):

wait so what is the actual answer, D?

OpenStudy (anonymous):

@shamil98

OpenStudy (shamil98):

are you serious..

OpenStudy (the_fizicx99):

LOL

OpenStudy (anonymous):

sorry...

OpenStudy (charlotte123):

No no - THE ANSWER IS G! \(\huge\color{black}{\frak{For~G~Is~Always~The~Answer}}\)

OpenStudy (charlotte123):

@itsannalea I saw that - No need to hide troubled mortal.

OpenStudy (anonymous):

wait what? I'm really confused I came on here for help and everyone is being weird.

OpenStudy (anonymous):

I deleted 'wth' because it was irrelevant.

OpenStudy (anonymous):

@ChrisTopher987654321

OpenStudy (charlotte123):

Nothing here Is weird my fellow - You need to understand Sarcasm.

OpenStudy (anonymous):

...

OpenStudy (the_fizicx99):

U_U

OpenStudy (shamil98):

i'm still wondering why you can't determine the answer. 0.05n + 0.10(n-6) = 2.10 like srsly.

OpenStudy (shamil98):

:(

OpenStudy (the_fizicx99):

@itsannalea he gave you the answer ._. th

OpenStudy (anonymous):

I have a major gap in math because earlier this year my best friend died so we were out of the state (therefore out of school) for a while.

OpenStudy (anonymous):

@tHe_FiZiCx99 this has nothing to do with you.

OpenStudy (the_fizicx99):

He just changed the format, thats all. Look it over and you'll have your answer ._.

OpenStudy (charlotte123):

Today Is going to be a really messed up Day -

OpenStudy (charlotte123):

Like - For real

OpenStudy (shamil98):

0.05n + 0.10(n-6) = 2.10 is the same as 0.10(n-6) + 0.05n = 2.10 when doing addition and multiplication the order does not matter. this is known as the associative property @itsannalea

OpenStudy (shamil98):

does that clear it up?

OpenStudy (anonymous):

not really but okay.

OpenStudy (shamil98):

that is a basic arithmetic rule, i expected you to know this since you are doing algebra.

OpenStudy (anonymous):

Well, guess what...

OpenStudy (anonymous):

guys honestly wth like I didn't do anything...

OpenStudy (charlotte123):

Neither did we Sir - Neither did we

OpenStudy (shamil98):

http://www.mathsisfun.com/associative-commutative-distributive.html read that if you need more of an understanding on that rule.

OpenStudy (anonymous):

I'm obviously trying to get help I'm not too good with math and I'm shocked that I'm even in Algebra 1.

OpenStudy (anonymous):

I never said that I understand it or anything...

OpenStudy (the_fizicx99):

What is that you need clarification on :o

OpenStudy (anonymous):

@shamil98 I UNDERSTAND THE RULE I just wanted to know what the actual answer was-therefore I put the choices in.

OpenStudy (anonymous):

@charlotte123 BTW I'm not a 'sir'

OpenStudy (charlotte123):

@itsannalea Sorry madam.